Hartman Motors has $10 million in assets, which were financed with $4 million of debt and $6 million in equity. Hartman's beta is currently 1.7, and its tax rate is 40%. Use the Hamada equation to find Hartman's unlevered beta, bU. Do not round intermediate calculations. Round your answer to two decimal places.

bU =
Answer #1

Unlevered Beta = Levered beta /[ (1+ (1-Tax rate ) (Debt / equity)]

Unlevered Beta = 1.7 /[(1+(1-0.4) (4/6)]

Unlevered Beta = 1.7 /(1 + (0.6) (0.66667)= 1.7/ 1.4= 1.214 or 1.21